OpenStudy (anonymous):

The width of a rectangle is half the length. The perimeter of the rectangle is 54 in. What is the length of the rectangle? a. 6 in. c. 18 in. b. 9 in. d. 36 in.

OpenStudy (ivettef365):

P = 2L + 2W

OpenStudy (ivettef365):

then you have P = 54 L = 2L W = 2 (1/2L)

OpenStudy (ivettef365):

so you solve it like this 2L +2(1/2L) = 54 solve for L

OpenStudy (anonymous):

bleh i still dunno

OpenStudy (ivettef365):

ok, that is the same as 2L + L = 54 just because 2(1/2L) = 1L can you take it from there ?

OpenStudy (anonymous):

mhm ok thnx

OpenStudy (ivettef365):

Do you know how to solve for L ??

OpenStudy (anonymous):

mhmm now i do

OpenStudy (ivettef365):

just add 2L + L = 54 then you have 3L = 54 , then divide by 3, so what do you get ??

OpenStudy (anonymous):

i had got 18

OpenStudy (ivettef365):

good job

OpenStudy (anonymous):

ty :P